OpenStudy (anonymous):

In Ron's class, the ratio of students who have pets to the students who do not have pets is 6:8. Which is an equivalent ratio?

OpenStudy (anonymous):

6 to8 =6:8=6/8= 3/4

OpenStudy (anonymous):

where is the choices

OpenStudy (anonymous):

2:3 8:12 12:16 24:40 these are the choices

jigglypuff314 (jigglypuff314):

it's like saying 6 -- 8 what is an equivalent fraction?

OpenStudy (anonymous):